Perimeter of a triangle is 69, side a is 5 inches shorter then side b, Side c is twice as long as side a. Find the length of each side
Answer by nyc_function(2741)

I will give you the equation and you can then solve for each side.

side a = (x - 5)

side b = x

side c = 2(x - 5)

Perimeter is 69.

69 = (x - 5) + x + 2(x - 5)

Can you finish?